OpenStudy (anonymous):

for the following set of data find an appropriate regression equation (linear, exponential, or logarithmic) state your reason x: 1,2,3,4,5 y: 112,114, 119, 122, 129

OpenStudy (anonymous):

\[\sum_{n=1}^{3}i^(n-1)\]

OpenStudy (anonymous):

in the second equation the right side of the sigma should read i^(n-1)

OpenStudy (bahrom7893):

do you mean i as in imaginary number?

OpenStudy (bahrom7893):

\[\sum_{n=1}^{3}i^{(n-1)}\] where i^2 =-1?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

\[\sum_{n=1}^{3}i^{n-1} = i^{1-1} + i^{2-1} + i^{3-1}\]

OpenStudy (bahrom7893):

\[i^0 + i + i^2 = 1 + i -1 = i\]

OpenStudy (anonymous):

okay thats what i got, thank you. can you take a look at the problem above that

OpenStudy (bahrom7893):

oh for that u need to use a calculator, do u want me to write the equation of the line?

OpenStudy (anonymous):

yeah i understand that but how would you know which type of regression to use?

OpenStudy (bahrom7893):

I think you just try every single one, and the one that goes through the most points will be the one to use, because calculator would approximate all equations. So just try all three, and whichever one comes closest to as many points as possible will be the one to use.

OpenStudy (bahrom7893):

well, that's how I used to do it, back in hs

OpenStudy (anonymous):

how do i graph it? i put in the type of regression and then (L1,L2) then how do I graph it?

OpenStudy (bahrom7893):

okay what calculator are you using?

OpenStudy (bahrom7893):

Ti 84? 89? 83?

OpenStudy (anonymous):

TI-83

OpenStudy (anonymous):

yeah i got that but my graph axis are bolded i dont know how to get out of that i clicked zoom out a couple of times but it wont work

OpenStudy (bahrom7893):

try zoom fit

OpenStudy (anonymous):

its not doing antyhing

OpenStudy (bahrom7893):

try decimal

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

i got it thank you

OpenStudy (bahrom7893):

okay i gtg now and that meant zoom decimal

OpenStudy (anonymous):

nothings graphing though

OpenStudy (bahrom7893):

which means u have to try a different regression

OpenStudy (bahrom7893):

and make sure the plots are selected in the y-editor

OpenStudy (bahrom7893):

sorry i gtg now, will try to respond later again

OpenStudy (anonymous):

yeah they are

OpenStudy (anonymous):

ok thanks